OpenStudy (anonymous):

The sum of two consecutive integers is -201. Find the larger ?

OpenStudy (anonymous):

x+ (x+1)=-201

OpenStudy (anonymous):

2x=-202

OpenStudy (anonymous):

x=-101

OpenStudy (anonymous):

so -100-101=-201

OpenStudy (anonymous):

x+ (x-1)=-201

OpenStudy (anonymous):

that is -101 and -101

OpenStudy (anonymous):

2x=-200

OpenStudy (anonymous):

x=-100

OpenStudy (anonymous):

so -100 and -101

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

-100

OpenStudy (anonymous):

meaning the larger is -100

OpenStudy (anonymous):

yes